OpenStudy (anonymous):

What is 1,430,000,000 in scientific notation?

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what i think

OpenStudy (madhu.mukherjee.946):

1.43*10^9

OpenStudy (anonymous):

but its only 7 zeros!

OpenStudy (madhu.mukherjee.946):

calculate it in the calculator

OpenStudy (anonymous):

lol ok

OpenStudy (madhu.mukherjee.946):

there are 7 zeros but there is two decimal places ,1.43

hartnn (hartnn):

7 zeros so its \(143 \times 10^7\)

hartnn (hartnn):

which is \(14.3 \times 10^8\) which is \(1.43 \times 10^9\)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

thanks!!

hartnn (hartnn):

welcome ^_^